OpenStudy (anonymous):

One thousand dollars is invested in a savings account that pays 9% interest per year. The interest earned after the first year is added to the account. How much interest is earned on the new principal the following year?

OpenStudy (kropot72):

9% of $1000 = $90 Amount at start of second year = 1000 + 90 = $1090 9% of $1090 = 1090 * 0.09 = $98.10 Interest on new principal = $98.10